If x+y+z=15 when a,x,y,z,b are in AP and 1/x+1/y+1/z=5/3 when a,x,y,z,b are in HP then


a.GM of a & b is 3


b.AM of a & b is 6


c.1 possible value of a+2b is 11


d.greatest value of a-b is 8

Grade:

2 Answers

Rup Chitrak
12 years ago

[Ans: (a),(c),(d)]

for AP, let x=a+d, y=a+2d, z=a+3d.

x+y+z=15

or, 3a+6d=15

or, a+2d=5 or, y=5

but y=b-2d

2y=10

a+b=10 ...(i)

for HP, let 1/x = 1/a + f ,  1/y = 1/a + 2f ,     1/z = 1/a + 3f

1/x + 1/y + 1/z = 5/3

or, 3/a + 6f = 5/3  or,  1/a + 2f = 5/9.

or, 1/y =5/9 or, 1/b - 2f = 5/9.

thus, 1/a + 1/b = 10/9  ....(ii)

from (i) and (ii)

ab = 9  or, (ab)1/2 = 3

solve (i) and (ii) to get other two options.Cool

Arkajyoti Batabyal
6 years ago
The system is inconsistent as y=5 (x,y,z are in AP) and y=9/5 (x,y,z are in HP) which is not possible.Thanks